A U-tube contains two fluids with densities ρ1 = 1065 kg/m3 and ρ2 = 540 kg/m3 as sketched below. What is the difference d in the heights of the top surfaces of the two fluids?
Answer #1

here figure is not given

but i can give some hint regarding this problem

at the bottom pressure nust same for both the fluids

p1= p2

p0+rho1gh1 = p0+rho2gh2

rho1h1 = rho2h2

h1/h2 = rho2/rho1 = 540/1065 = 0.507